OpenStudy (anonymous):

find y' assuming that the equation deermines a differentiable function f such that y =f(x) y^2+1=x^2 3y topic: implicit differentiation

OpenStudy (anonymous):

y^2 +1 = x^2 sec y

OpenStudy (amistre64):

3y or secy?

OpenStudy (anonymous):

take the derivative of both sides wrt x

OpenStudy (anonymous):

2yy'=2x sec(y) +x^2 sec(y)tan(y)y'

OpenStudy (anonymous):

solve for y' using algebra

OpenStudy (anonymous):

amistre hello!

OpenStudy (anonymous):

i think it was \[y^2=x^2\sec(y)\] yes?

OpenStudy (amistre64):

howdy.... openstudys behaving bad tonight. i go to post and i get screenfreeze

OpenStudy (anonymous):

me too!

OpenStudy (amistre64):

2y y' = 2x3y + x^2 3 y' 2y y' -x^2 3 y' = 6xy y'(2y -3x^2) = 6xy y' = 6xy/(2y-3x^2)

OpenStudy (anonymous):

y^2 + 1 = x^2 sec y satellite

OpenStudy (anonymous):

so i used the product rule for the x^2 sec y part?

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

can you guys help me with the question i just posted

OpenStudy (anonymous):

yes the product rule. and when you take the derivative of sec(y) make sure to put the y' at the end. i will write it out if you like let me kn ow

OpenStudy (anonymous):

i can do it thank you

OpenStudy (anonymous):

2yy' = (2x)(secy)+(secy'tany')(x^2) is it correct?